OpenStudy (anonymous):

4xy+3=5z Find The Value Of Y ? 2m-3p=1 Find the Value Of P a+b=2xz Find The Value Of Z ?

OpenStudy (anonymous):

can you help me with these ones too @grifter192

OpenStudy (anonymous):

4xy+3=5z 4xy = 5z - 3 y = (5z - 3)/4x 2m-3p=1 2m = 1 + 3p 2m - 1 = 3p (2m - 1)/3 = p a+b=2xz (a+b)/2x = z